This year, FCF, Inc. Has earnings before interest and taxes of $10 million, depreciation expenses of $1 million, capital expenditures of $1.5 million and has increased its net working capital by $500,000. If its tax rate is 35%, what is its free cash flow?
Answer #1

Based on mathematical relationship,

FCF = EBIT * (1 - Tax Rate) + Depreciation - Changes in WC - Capex

FCF = 10 * (1 - 35%) + 1 - 0.5 - 1.5

FCF = 6.5 + 1 - 2

FCF = $5.5 mil

FCF = $5,500,000